OpenStudy (anonymous):

The zero of F-1(x) in F(x) = x + 3 is

OpenStudy (amoodarya):

y=x+3 x=y-3 so f-1(x)=x-3

OpenStudy (anonymous):

but I need the zero

OpenStudy (anonymous):

How do you find the zero?

OpenStudy (amoodarya):

ok lets do that f-1(x)=0 x-3=0 then x=3

OpenStudy (anonymous):

Oh okay that's understandable thanks, I was really confused!